OpenStudy (anonymous):

Solve for C: F=9/5C+32

OpenStudy (marigirl):

lets subtract 32 from both sides. F-32=(9/5)C then multiply both sides by 5 5(F-32)=9c then divide both sides by 9 C= 5(F-32)/9

OpenStudy (marigirl):

can u see attached?

OpenStudy (anonymous):

Thank u sooo much!(;

OpenStudy (marigirl):

cool, hope u were able to see my attached file cuz equation is not working on here for me :(